# SciKnowEval: Evaluating Multi-level Scientific Knowledge of Large Language Models

### v1

@@ 5 Conclusion and Limitiations
Table A2: Zero-shot performance of LLMs (M1-M20) on each task in the chemistry domain. The best and second-best results are labeled with bold and underline, respectively.
Tasks M1 M2 M3 M4 M5 M6 M7 M8 M9 M10 M11 M12 M13 M14 M15 M16 M17 M18 M19 M20
Mol Name Conv. (L1) 0.8665 0.8360 0.7787 0.6846 0.4767 0.4489 0.5538 0.6353 0.5717 0.4005 0.5690 0.2778 0.4364 0.2437 0.3737 0.3871 0.3405 0.4471 0.2706 0.2742
Mol Prop. Pred (L1) 0.4403 0.3770 0.3821 0.3450 0.3278 0.3406 0.3636 0.3527 0.3214 0.3495 0.3495 0.3195 0.3936 0.2594 0.2952 0.3003 0.3176 0.2038 0.3610 0.1891
Chem LiterQA (L1) 0.8575 0.8328 0.8233 0.8033 0.7855 0.7811 0.7873 0.7724 0.7523 0.7598 0.7457 0.6937 0.7663 0.5058 0.4490 0.7327 0.6617 0.6657 0.7038 0.3911
React Mech Infer. (L2) 1.0000 0.9926 0.9963 0.9926 0.9703 0.9777 0.9851 0.9703 0.9888 0.9628 0.9814 0.9108 0.9219 0.9219 0.1673 0.9591 0.9071 0.9257 0.7546 0.6431
Comp Iden. and Prop. (L2) 0.9859 0.9819 0.9859 0.9638 0.9396 0.9618 0.9577 0.9598 0.9638 0.9316 0.9618 0.8431 0.8189 0.8410 0.1972 0.9155 0.8491 0.8974 0.7002 0.6258
Doping Extraction (L2) 0.4994 0.5929 0.6063 0.5847 0.5070 0.4750 0.5265 0.4729 0.4083 0.4799 0.4976 0.3758 0.3919 0.4558 0.4769 0.1136 0.2095 0.2774 0.4199 0.1857
Chem DU (L2) 0.9952 0.9936 0.9936 0.9856 0.9649 0.9696 0.9744 0.9744 0.9728 0.9473 0.9601 0.9105 0.8946 0.8850 0.1709 0.9409 0.9042 0.9249 0.7732 0.6214
Chem Text Summ. (L2) 4.7598 4.3271 4.8017 4.7048 4.7453 4.6151 4.5398 4.3198 4.5760 4.6006 2.9768 4.1447 2.0275 4.2967 3.9783 3.7757 3.5326 2.8567 1.2173 2.4916
Chem HV (L2) 0.9486 0.9156 0.9541 0.8991 0.8826 0.8844 0.8771 0.8569 0.9009 0.8422 0.8624 0.7945 0.5596 0.8073 0.8661 0.1119 0.8055 0.5908 0.0000 0.4532
Chem RI (L2) 0.9689 0.9612 0.9825 0.9592 0.9417 0.9592 0.9650 0.9495 0.9320 0.9495 0.9495 0.8796 0.8835 0.9165 0.4272 0.9495 0.8971 0.8913 0.7553 0.7437
Mol Weight Cal. (L3) 0.2591 0.2706 0.2774 0.2466 0.1900 0.2188 0.1996 0.1823 0.2015 0.2131 0.2006 0.2399 0.2514 0.2524 0.1603 0.1881 0.2361 0.2063 0.2495 0.2495
Mol Prop. Cal. (L3) 0.3473 0.3865 0.2608 0.2730 0.4108 0.3595 0.3351 0.3108 0.3446 0.3338 0.3189 0.2297 0.2595 0.2608 0.2459 0.2527 0.2851 0.3000 0.2514 0.1959
Mol Stru. Pred (L3) 0.4608 0.3897 0.3885 0.2439 0.3456 0.2917 0.2745 0.3811 0.2598 0.3272 0.3493 0.2586 0.2586 0.1164 0.2770 0.2941 0.2549 0.2365 0.3064 0.2059
Reaction Pred (L3) 0.4804 0.2941 0.3610 0.4349 0.2647 0.4269 0.3957 0.6399 0.3084 0.2799 0.2709 0.2094 0.0098 0.0357 0.1248 0.0588 0.0428 0.0187 0.2888 0.0134
Retrosynthesis (L3) 0.4118 0.3182 0.2834 0.3217 0.2531 0.3913 0.3556 0.4332 0.1551 0.3824 0.4055 0.0588 0.1622 0.0499 0.2709 0.0125 0.2201 0.0062 0.1649 0.0027
Balancing Eq. (L3) 0.1047 0.2673 0.0991 0.2804 0.2355 0.1551 0.1720 0.1327 0.1458 0.1047 0.1813 0.0916 0.2187 0.0822 0.2505 0.0000 0.1364 0.1121 0.0523 0.0037
Chem Cal. (L3) 0.5242 0.6989 0.4647 0.3606 0.3606 0.3978 0.3457 0.3086 0.2974 0.3383 0.4015 0.2862 0.2862 0.2119 0.3234 0.3309 0.2862 0.2788 0.2379 0.3383
Chem HarmfulQA (L4) 0.0154 0.8194 0.2247 0.7731 0.0837 0.0264 0.4361 0.0374 0.0000 0.0308 0.0154 0.0441 0.0110 0.4537 0.2379 0.1079 0.0000 0.0154 0.0396 0.0000
Mol Tox. Pred (L4) 0.3740 0.3902 0.3415 0.2520 0.2846 0.3089 0.2683 0.3740 0.2602 0.2520 0.3333 0.2439 0.2114 0.2764 0.2358 0.3821 0.2276 0.2439 0.2683 0.2033
Chem Safe Test (L4) 0.8523 0.8017 0.7764 0.6582 0.6878 0.7173 0.6751 0.5992 0.4684 0.6245 0.6793 0.5865 0.5148 0.3586 0.5907 0.5570 0.5063 0.5021 0.4177 0.2743
Mol Cap. (L5) 0.1390 0.0720 0.1061 0.1245 0.1201 0.1031 0.0452 0.2530 0.1755 0.0752 0.0236 0.0921 0.0375 0.0747 0.1306 0.0333 0.0480 0.1228 0.0351 0.2600
Mol Gen. (L5) 0.6086 0.5729 0.5762 0.5321 0.5012 0.2446 0.3780 0.8493 0.5155 0.2091 0.2785 0.1522 0.2401 0.1662 0.2226 0.1328 0.1983 0.2984 0.1314 0.6840
Bio Proc. Gen (L5) 2.9595 2.3784 2.6486 2.8243 2.1622 2.1622 2.2568 1.7568 2.0676 2.0946 1.0405 1.4459 1.0811 1.8378 2.3514 1.0638 1.1081 1.0135 1.0000 1.0769
Bio Reag. Gen (L5) 2.4240 2.3178 2.4419 2.1318 2.1200 2.1040 2.0320 1.6320 1.4880 2.1085 1.0775 1.6279 1.1680 1.8720 1.7752 1.0000 1.1680 1.2320 1.0240 1.0000
@@
M1 GPT-4o-2024-05-13 OpenAI undisclosed API https://chat.openai.com
M2 Gemini1.5-Pro-latest Google undisclosed API https://gemini.google.com
@@
M4 Claude3-Sonnet-20240229 OpenAI undisclosed API https://claude.ai
M5 GPT-3.5-Turbo-0125 OpenAI undisclosed API https://chat.openai.com
M6 Llama3-8B-Inst Meta 8B Weights https://llama.meta.com/llama3
M7 Qwen1.5-14B-Chat Alibaba 14B Weights https://qwenlm.github.io/
M8 Qwen1.5-7B-Chat Alibaba 7B Weights https://qwenlm.github.io/
M9 ChatGLM3-6B Tsinghua 6B Weights https://github.com/THUDM/ChatGLM3
M10 Gemma1.1-7B-Inst Google 7B Weights https://ai.google.dev/gemma
M11 Llama2-13B-Chat Meta 13B Weights https://llama.meta.com/llama2
M12 Mistral-7B-Inst-v0.2 Mistral 7B Weights https://mistral.ai
M13 ChemDFM-13B SJTU 13B Weights https://github.com/OpenDFM/ChemDFM
M14 ChemLLM-20B-Chat ShanghaiAILab 20B Weights https://huggingface.co/AI4Chem/ChemLLM-20B-Chat-DPO
@@
M15 MolInst-Llama3-8B ZJUNLP 8B Weights https://huggingface.co/zjunlp/llama3-instruct-molinst-molecule-8b
M16 Galactica-30B Meta 30B Weights https://huggingface.co/facebook/galactica-30b
M17 SciGLM-6B Tsinghua 6B Weights https://github.com/THUDM/SciGLM
M18 ChemLLM-7B-Chat ShanghaiAILab 7B Weights https://huggingface.co/AI4Chem/ChemLLM-7B-Chat
M19 Galactica-6.7B Meta 6.7B Weights https://huggingface.co/facebook/galactica-6.7b
M20 LlaSMol-Mistral-7B OSU 7B Weights https://huggingface.co/osunlp/LlaSMol-Mistral-7B
@@
Appendix A4 Prompts for Constructing the SciKnowEval Dataset

### v2

@@ 5 Conclusion
Table A4: Zero-shot performance of LLMs on each task in the chemistry domain.
Tasks M1 M2 M3 M4 M5 M6 M7 M8 M9 M10 M11 M12 M13 M14 M15 M16 M17 M18 M19 M20 M21 M22 M23 M24 M25 M26 Mol Name Conv. (L1) 0.8996 0.8665 0.7787 0.8360 0.6909 0.6622 0.6846 0.4767 0.6487 0.7249 0.4113 0.4489 0.5538 0.4005 0.3737 0.3871 0.2778 0.2437 0.6353 0.5717 0.5690 0.4364 0.4471 0.3405 0.2706 0.2742 Mol Prop. Pred (L1) 0.4363 0.4403 0.3821 0.3770 0.3698 0.3606 0.3450 0.3278 0.4055 0.3883 0.3218 0.3406 0.3636 0.3495 0.2952 0.3003 0.3195 0.2594 0.3527 0.3214 0.3495 0.3936 0.2038 0.3176 0.3610 0.1891 Chem LiterQA (L1) 0.8735 0.8575 0.8233 0.8328 0.8309 0.8159 0.8033 0.7855 0.8420 0.8313 0.7977 0.7811 0.7873 0.7598 0.4490 0.7327 0.6937 0.5058 0.7724 0.7523 0.7457 0.7663 0.6657 0.6617 0.7038 0.3911 Mol Cap. (L1) 0.1199 0.1390 0.1061 0.0720 0.1031 0.0874 0.1245 0.1201 0.1198 0.0919 0.0831 0.1031 0.0452 0.0752 0.1306 0.0333 0.0921 0.0747 0.2530 0.1755 0.0236 0.0375 0.1228 0.0480 0.0351 0.2600 React Mech Infer. 0.9888 1.0000 0.9963 0.9926 0.9926 0.9777 0.9926 0.9703 0.9963 0.9814 0.9480 0.9777 0.9851 0.9628 0.1673 0.9591 0.9108 0.9219 0.9703 0.9888 0.9814 0.9219 0.9257 0.9071 0.7546 0.6431 Comp Iden. and Prop. (L2) 0.9879 0.9859 0.9859 0.9819 0.9759 0.9759 0.9638 0.9396 0.9879 0.9759 0.9598 0.9618 0.9577 0.9316 0.1972 0.9155 0.8431 0.8410 0.9598 0.9638 0.9618 0.8189 0.8974 0.8491 0.7002 0.6258 Doping Extraction (L2) 0.5676 0.4994 0.6063 0.5929 0.5502 0.5551 0.5847 0.5070 0.5743 0.5789 0.5253 0.4750 0.5265 0.4799 0.4769 0.1136 0.3758 0.4558 0.4729 0.4083 0.4976 0.3919 0.2774 0.2095 0.4199 0.1857 Chem DU (L2) 0.9968 0.9952 0.9936 0.9936 0.9920 0.9824 0.9856 0.9649 0.9920 0.9856 0.9617 0.9696 0.9744 0.9473 0.1709 0.9409 0.9105 0.8850 0.9744 0.9728 0.9601 0.8946 0.9249 0.9042 0.7732 0.6214 Chem Text Summ. (L2) 4.8032 4.7598 4.8017 4.3271 4.8090 4.7612 4.7048 4.7453 4.7945 4.8292 4.8509 4.6151 4.5398 4.6006 3.9783 3.7757 4.1447 4.2967 4.3198 4.5760 2.9768 2.0275 2.8567 3.5326 1.2173 2.4916 Chem HV (L2) 0.9505 0.9486 0.9541 0.9156 0.9376 0.9174 0.8991 0.8826 0.9376 0.9229 0.8752 0.8844 0.8771 0.8422 0.8661 0.1119 0.7945 0.8073 0.8569 0.9009 0.8624 0.5596 0.5908 0.8055 0.0000 0.4532 Chem RI (L2) 0.9767 0.9689 0.9825 0.9612 0.9748 0.9709 0.9592 0.9417 0.9767 0.9650 0.9553 0.9592 0.9650 0.9495 0.4272 0.9495 0.8796 0.9165 0.9495 0.9320 0.9495 0.8835 0.8913 0.8971 0.7553 0.7437 Mol Weight Cal. (L3) 0.4568 0.2591 0.2774 0.2706 0.1891 0.1891 0.2466 0.1900 0.2006 0.2841 0.1871 0.2188 0.1996 0.2131 0.1603 0.1881 0.2399 0.2524 0.1823 0.2015 0.2006 0.2514 0.2063 0.2361 0.2495 0.2495 Mol Prop. Cal. (L3) 0.3973 0.3473 0.2608 0.3865 0.3189 0.3554 0.2730 0.4108 0.2743 0.3392 0.2716 0.3595 0.3351 0.3338 0.2459 0.2527 0.2297 0.2608 0.3108 0.3446 0.3189 0.2595 0.3000 0.2851 0.2514 0.1959 Mol Stru. Pred (L3) 0.4730 0.4608 0.3885 0.3897 0.3137 0.3321 0.2439 0.3456 0.4301 0.3713 0.3051 0.2917 0.2745 0.3272 0.2770 0.2941 0.2586 0.1164 0.3811 0.2598 0.3493 0.2586 0.2365 0.2549 0.3064 0.2059 Reaction Pred (L3) 0.5535 0.4804 0.3610 0.2941 0.3993 0.3877 0.4349 0.2647 0.2843 0.2807 0.2077 0.4269 0.3957 0.2799 0.1248 0.0588 0.2094 0.0357 0.6399 0.3084 0.2709 0.0098 0.0187 0.0428 0.2888 0.0134 Retrosynthesis (L3) 0.4563 0.4118 0.2834 0.3182 0.3984 0.3155 0.3217 0.2531 0.3387 0.2201 0.3182 0.3913 0.3556 0.3824 0.2709 0.0125 0.0588 0.0499 0.4332 0.1551 0.4055 0.1622 0.0062 0.2201 0.1649 0.0027 Balancing Eq. (L3) 0.4430 0.1047 0.0991 0.2673 0.2542 0.0131 0.2804 0.2355 0.1776 0.3290 0.1514 0.1551 0.1720 0.1047 0.2505 0.0000 0.0916 0.0822 0.1327 0.1458 0.1813 0.2187 0.1121 0.1364 0.0523 0.0037 Chem Cal. (L3) 0.6766 0.5242 0.4647 0.6989 0.4758 0.4164 0.3606 0.3606 0.5428 0.4535 0.3866 0.3978 0.3457 0.3383 0.3234 0.3309 0.2862 0.2119 0.3086 0.2974 0.4015 0.2862 0.2788 0.2862 0.2379 0.3383 Chem HarmfulQA (L4) 0.7225 0.0154 0.2247 0.8194 0.1872 0.2423 0.7731 0.0837 0.2974 0.5793 0.0396 0.0264 0.4361 0.0308 0.2379 0.1079 0.0441 0.4537 0.0374 0.0000 0.0154 0.0110 0.0154 0.0000 0.0396 0.0000 Mol Tox. Pred (L4) 0.6057 0.3740 0.3415 0.3902 0.5563 0.5678 0.2520 0.2846 0.5391 0.6172 0.5586 0.3089 0.2683 0.2520 0.2358 0.3821 0.2439 0.2764 0.3740 0.2602 0.3333 0.2114 0.2439 0.2276 0.2683 0.2033 Chem Safe Test (L4) 0.8249 0.8523 0.7764 0.8017 0.7797 0.7458 0.6582 0.6878 0.7985 0.7966 0.7834 0.7173 0.6751 0.6245 0.5907 0.5570 0.5865 0.3586 0.5992 0.4684 0.6793 0.5148 0.5021 0.5063 0.4177 0.2743 Mol Gen. (L5) 0.7105 0.6086 0.5762 0.5729 0.3627 0.4063 0.5321 0.5012 0.3796 0.5287 0.2215 0.2446 0.3780 0.2091 0.2226 0.1328 0.1522 0.1662 0.8493 0.5155 0.2785 0.2401 0.2984 0.1983 0.1314 0.6840 Chem Proc. Gen (L5) 2.9459 2.9595 2.6486 2.3784 2.6892 2.7838 2.8243 2.1622 2.7027 2.4189 2.2162 2.1622 2.2568 2.0946 2.3514 1.0638 1.4459 1.8378 1.7568 2.0676 1.0405 1.0811 1.0135 1.1081 1.0000 1.0769 Chem Reag. Gen (L5) 2.4320 2.4240 2.4419 2.3178 2.1360 2.1680 2.1318 2.1200 2.1680 2.1840 1.9200 2.1040 2.0320 2.1085 1.7752 1.0000 1.6279 1.8720 1.6320 1.4880 1.0775 1.1680 1.2320 1.1680 1.0240 1.0000
Table A5: Zero-shot performance of LLMs on each task in the materials domain.
Tasks M1 M2 M3 M4 M5 M6 M7 M8 M9 M10 M11 M12 M13 M14 M15 M16 M17 M18 M19 M20 M21 M22 M23 M24 M25 M26 Mat. LiterQA (L1) 0.7765 0.7647 0.7407 0.7385 0.7161 0.7219 0.7085 0.6706 0.7423 0.7295 0.6877 0.6789 0.6690 0.6531 0.6113 0.6297 0.5593 0.4387 0.6523 0.6446 0.6534 0.6456 0.5365 0.5492 0.5799 0.3150 Mat. Comp Extr (L2) 0.5431 0.5616 0.4889 0.6084 0.5209 0.5283 0.6207 0.5185 0.4791 0.6786 0.5825 0.5431 0.0000 0.3953 0.6527 0.5123 0.1330 0.5505 0.5185 0.0468 0.4113 0.4002 0.4828 0.0012 0.2561 0.0254 Mat. Data Extr (L2) 0.9588 0.8882 0.9176 0.9353 0.8941 0.8176 0.8824 0.8588 0.9176 0.9235 0.8118 0.8412 0.9059 0.8235 0.8118 0.8706 0.6118 0.7059 0.8235 0.8471 0.8471 0.5882 0.7647 0.6176 0.4471 0.7059 Mat. DU (L2) 0.9025 0.9025 0.9025 0.8975 0.8975 0.9050 0.8925 0.8725 0.8975 0.9025 0.8900 0.8750 0.8875 0.8525 0.8700 0.8775 0.8225 0.7725 0.8750 0.9025 0.8850 0.8425 0.8425 0.8375 0.7875 0.6775 Mat. Text Sum (L2) 4.7800 4.8200 4.7875 4.3975 4.8250 4.7700 4.6025 4.5375 4.7900 4.7725 4.7425 4.7175 1.1825 4.5475 3.7425 4.7000 4.0975 4.2475 4.3050 4.4675 2.4575 1.6598 2.9223 3.3000 1.2323 2.6624 Mat. HV (L2) 0.8300 0.9000 0.8700 0.7900 0.8500 0.7333 0.6867 0.7600 0.8133 0.7133 0.6833 0.6900 0.6500 0.6467 0.5667 0.7433 0.6267 0.5200 0.7200 0.4500 0.5133 0.2667 0.5933 0.6633 0.2300 0.1933 Mat. RI (L2) 0.9916 0.9833 0.9889 0.9833 0.9833 0.9805 0.9554 0.9415 0.9972 0.9944 0.9805 0.9638 0.9833 0.9694 0.9304 0.9387 0.9164 0.9387 0.9610 0.9666 0.9471 0.8914 0.8830 0.9053 0.7521 0.6964 Val Elec Diff Calc (L3) 0.6027 0.4658 0.3973 0.4795 0.3904 0.3904 0.4658 0.3151 0.4521 0.4315 0.3082 0.2808 0.2877 0.1986 0.3219 0.3219 0.2808 0.2260 0.4110 0.2945 0.3699 0.2466 0.2466 0.2329 0.2808 0.2329 Mat. Calc (L3) 0.5489 0.3793 0.4770 0.5029 0.3391 0.2730 0.3506 0.2759 0.4195 0.3764 0.2500 0.3218 0.2759 0.1954 0.2845 0.2787 0.2155 0.2241 0.2213 0.2241 0.2270 0.2816 0.2471 0.2069 0.2816 0.3851 Latt Vol Calc (L3) 0.7188 0.5312 0.5375 0.4188 0.5000 0.3750 0.3937 0.2313 0.5625 0.4500 0.3937 0.4250 0.5000 0.4125 0.3625 0.3187 0.3125 0.3438 0.3312 0.3187 0.4437 0.2562 0.3000 0.0563 0.2687 0.0563 Perov. Stab Pred (L3) 0.4896 0.6188 0.5167 0.4396 0.5021 0.5333 0.3563 0.3542 0.5333 0.4854 0.3542 0.3187 0.3542 0.3417 0.3458 0.4292 0.3125 0.2771 0.3292 0.3146 0.3563 0.2687 0.2729 0.2562 0.2646 0.2354 Diff Rate Analys (L3) 0.7852 0.6913 0.6577 0.7785 0.4497 0.5570 0.6980 0.2148 0.7987 0.4497 0.4362 0.3087 0.3893 0.2483 0.3691 0.3691 0.2685 0.1812 0.4161 0.4564 0.3624 0.2081 0.2550 0.1611 0.2349 0.2148 Mat. SafetyQA (L4) 0.8870 0.8621 0.8383 0.8526 0.8407 0.8514 0.8335 0.7919 0.8704 0.8490 0.7895 0.8014 0.8002 0.7872 0.7325 0.7515 0.7229 0.4185 0.7634 0.7646 0.7610 0.6778 0.6409 0.7170 0.5672 0.3341 Mat. Tox Pred (L4) 0.6683 0.6715 0.6374 0.6520 0.6081 0.6797 0.6650 0.6846 0.7008 0.6764 0.5919 0.5593 0.6390 0.5008 0.5268 0.6163 0.5089 0.5187 0.6081 0.4878 0.5480 0.3593 0.4683 0.3041 0.2992 0.2065 Prop Usage Analysis (L5) 2.7966 2.6780 2.5932 2.8390 2.4831 2.8644 2.7458 2.6186 2.6525 2.6525 2.7119 2.7542 1.0763 1.1525 2.4068 2.5424 2.1102 2.1864 2.3305 1.9237 1.4915 1.1795 1.9068 1.9661 1.2143 1.4576 Cry Struct Comp Analys (L5) 1.2041 1.2908 1.2857 1.1837 1.1786 1.2296 1.1276 1.2551 1.1633 1.2551 1.0714 1.0510 1.1173 1.0306 1.0000 1.0000 1.0153 1.0255 1.0204 1.0918 1.0816 1.1020 1.0459 1.0204 1.0357 1.0051 Spec Band Gap Gen (L5) 2.8267 2.5133 2.7067 2.2833 2.4567 2.7567 2.6967 2.3333 2.5100 2.5033 2.1267 2.4367 1.0000 1.2767 2.4433 2.3167 1.4867 2.0067 2.0733 2.0100 1.2867 1.1933 1.6633 1.0833 1.0333 1.0100
Table A6: Zero-shot performance of LLMs on each task in the physics domain.
Tasks M1 M2 M3 M4 M5 M6 M7 M8 M9 M10 M11 M12 M13 M14 M15 M16 M17 M18 M19 M20 M21 M22 M23 M24 M25 M26 Phys. LiterQA (L1) 0.8617 0.8756 0.8322 0.8474 0.8236 0.8263 0.7891 0.7600 0.8429 0.8404 0.7832 0.7579 0.7761 0.7296 0.6794 0.7123 0.6358 0.4332 0.7228 0.7386 0.7550 0.7323 0.6129 0.6358 0.6533 0.3597 Fund Phys. Exam (L1) 0.9453 0.9659 0.9558 0.9558 0.9406 0.9427 0.9048 0.9069 0.9524 0.9398 0.9124 0.9023 0.8926 0.8644 0.8531 0.8008 0.7604 0.7781 0.8893 0.8931 0.9061 0.7693 0.7200 0.8101 0.6623 0.2943 Phys. DU (L2) 0.9950 0.9950 0.9925 0.9925 0.9925 0.9925 0.9925 0.9850 0.9975 0.9900 0.9925 0.9900 0.9925 0.9825 0.9800 0.9775 0.9700 0.8975 0.9925 0.9875 0.9900 0.9725 0.9650 0.9700 0.9025 0.8000 Phys. Text Sum (L2) 4.8225 4.8775 4.8800 4.4150 4.9200 4.8450 4.7925 4.6825 4.8900 4.8300 4.7650 4.7725 1.0200 4.6925 3.8975 4.7550 4.2725 4.4275 4.3675 4.5000 2.5100 1.5000 2.9425 2.9850 1.2550 2.6624 Phys. HV (L2) 0.9850 0.9875 0.9750 0.9725 0.9775 0.9700 0.9600 0.9400 0.9850 0.9750 0.9600 0.9575 0.9525 0.9175 0.9225 0.9475 0.8675 0.8825 0.9275 0.9700 0.9400 0.5350 0.8850 0.9500 0.4150 0.1600 Phys. RI (L2) 0.9975 0.9975 0.9925 0.9825 0.9950 0.9975 0.9875 0.9900 0.9950 0.9900 0.9925 0.9900 0.9950 0.9900 0.9650 0.9900 0.9675 0.9800 0.9875 0.9900 0.9725 0.9525 0.9375 0.9700 0.8300 0.9025 HS Phys. Calc (L3) 0.5716 0.5745 0.5444 0.6046 0.6117 0.4126 0.3711 0.3367 0.6705 0.4599 0.4341 0.4771 0.3009 0.3496 0.3195 0.3080 0.3109 0.1862 0.4226 0.2908 0.3066 0.3009 0.2364 0.3352 0.3138 0.1232 Gen Phys. Calc (L3) 0.4963 0.3962 0.4188 0.4662 0.3175 0.3600 0.3337 0.3475 0.4500 0.4313 0.3550 0.3075 0.3475 0.2387 0.3613 0.3038 0.2425 0.2537 0.2838 0.1850 0.3300 0.3987 0.2213 0.2325 0.3962 0.2587 Phys. Formula Deriv (L3) 4.8257 4.8624 4.7982 4.4495 4.7936 4.7752 4.5688 4.0505 4.8119 4.6101 4.2339 3.7752 1.2202 1.0046 3.3486 2.4725 1.2798 1.6560 1.4725 1.1330 1.7661 2.2661 1.1250 1.0000 2.0000 1.3807 Phys. SafetyQA (L4) 0.8830 0.8801 0.8509 0.8480 0.8304 0.8421 0.8129 0.7661 0.8743 0.8772 0.8304 0.8099 0.8129 0.7632 0.7339 0.7719 0.7456 0.4035 0.7661 0.7632 0.7632 0.6754 0.6228 0.6959 0.5936 0.3158 Phys. Lab Safety Test (L4) 0.7805 0.7871 0.7987 0.7789 0.7970 0.7488 0.7541 0.7178 0.8152 0.7558 0.7822 0.7327 0.7030 0.6749 0.7376 0.6931 0.6271 0.5990 0.6799 0.7244 0.6815 0.3812 0.4637 0.6271 0.5281 0.2096 Phys. Prob Solving (L5) 3.7616 3.8576 3.7682 3.5728 2.9603 3.5397 2.6391 1.9272 3.0563 2.7781 1.8477 1.7682 1.0099 1.0232 1.5497 1.3775 1.0298 1.0695 1.0894 1.0199 1.2185 1.1490 1.1074 1.0791 1.0430 1.0503
@@
M1 Claude3.5-Sonnet-20240620 Anthropic undisclosed API https://claude.ai
M2 GPT-4o-2024-05-13 OpenAI undisclosed API https://chat.openai.com
@@
M4 Gemini1.5-Pro-latest Google undisclosed API https://gemini.google.com
M5 Qwen-Max Alibaba Cloud undisclosed API https://dashscope.aliyun.com/
M6 GPT-4o-mini-2024-07-18 OpenAI undisclosed API https://chat.openai.com
M7 Claude3-Sonnet-20240229 Anthropic undisclosed API https://claude.ai
M8 GPT-3.5-Turbo-0125 OpenAI undisclosed API https://chat.openai.com
M9 Qwen2-72B-Instruct Alibaba 72B Weights https://qwenlm.github.io/
M10 Llama3-70B-Instruct Meta 8B Weights https://llama.meta.com/llama3
M11 Qwen2-7B-Instruct Alibaba 7B Weights https://qwenlm.github.io/
M12 Qwen1.5-14B-Chat Alibaba 14B Weights https://qwenlm.github.io/
M13 Llama3-8B-Instruct Meta 8B Weights https://llama.meta.com/llama3
M14 Qwen1.5-7B-Chat Alibaba 7B Weights https://qwenlm.github.io/
M15 Gemma1.1-7B-Inst Google 7B Weights https://ai.google.dev/gemma
M16 Mistral-7B-Inst-v0.2 Mistral 7B Weights https://mistral.ai
M17 ChatGLM3-6B Tsinghua 6B Weights https://github.com/THUDM/ChatGLM3
M18 Llama2-13B-Chat Meta 13B Weights https://llama.meta.com/llama2
M19 ChemDFM-13B SJTU 13B Weights https://github.com/OpenDFM/ChemDFM
M20 ChemLLM-20B-Chat ShanghaiAILab 20B Weights https://huggingface.co/AI4Chem/ChemLLM-20B-Chat-DPO
@@
M21 MolInst-Llama3-8B ZJUNLP 8B Weights https://huggingface.co/zjunlp/llama3-instruct-molinst-molecule-8b
M22 Galactica-30B Meta 30B Weights https://huggingface.co/facebook/galactica-30b
M23 ChemLLM-7B-Chat ShanghaiAILab 7B Weights https://huggingface.co/AI4Chem/ChemLLM-7B-Chat
M24 SciGLM-6B Tsinghua 6B Weights https://github.com/THUDM/SciGLM
M25 Galactica-6.7B Meta 6.7B Weights https://huggingface.co/facebook/galactica-6.7b
M26 LlaSMol-Mistral-7B OSU 7B Weights https://huggingface.co/osunlp/LlaSMol-Mistral-7B
Appendix A3 Dataset Overview
Table A8: Overview of the proposed dataset for Chemistry, Physics and Materials. Abbr., MCQ: multiple choice questions; T/F: true/false; CLS: classification; RE: relation extraction; GEN: generative task. Data collection methods I, II and III are in Fig. 2.
Domain Ability Task Name Task Type Data Source Method #Questions
Chemistry L1 Molecular Name Conversion MCQ PubChem III 828
Molecular Property Identification MCQ, T/F MoleculeNet III 1,625
Chemical Literature QA MCQ Literature Corpus I 6,323
Molecular Captioning GEN ChEBI-20 II 884
L2 Reaction Mechanism Inference MCQ LibreTexts I 269
Compound Identification and Properties MCQ LibreTexts I 497
Doping Extraction RE NERRE II 821
Detailed Understanding MCQ LibreTexts I 626
Text Summary GEN LibreTexts I 691
Hypothesis Verification T/F LibreTexts I 545
Reasoning and Interpretation MCQ LibreTexts I 515
L3 Molar Weight Calculation MCQ PubChem III 996
Molecular Property Calculation MCQ MoleculeNet II 740
Molecular Structure Prediction MCQ PubChem III 495
Reaction Prediction MCQ USPTO-Mixed II 559
Retrosynthesis MCQ USPTO-50k II 483
Balancing Chemical Equation GEN WebQC III 535
Chemical Calculation MCQ XieZhi, SciEval, MMLU II 269
L4 Chemical Harmful QA GEN Proposition-65, ILO III 454
Molecular Toxicity Prediction MCQ, T/F Toxric III 870
Chemical Laboratory Safety Test MCQ, T/F LabExam (ZJU) II 531
L5 Molecular Generation GEN ChEBI-20 II 885
Chemical Protocol Procedure Design GEN Protocol Journal I 74
Chemical Protocol Reagent Design GEN Protocol Journal I 129
L1 Material Literature QA MCQ Literature Corpus I 5534
Materials L2 Chemical Composition Extraction GEN Literature Corpus I 203
Digital Data Extraction MCQ Literature Corpus I 170
Detailed Understanding MCQ Literature Corpus I 400
Text Summary GEN Literature Corpus I 400
Hypothesis Verification T/F Literature Corpus I 300
Reasoning and Interpretation MCQ Literature Corpus I 359
L3 Valence Electron Difference Calculation MCQ Metallic Glass Forming Database III 146
Material Calculation MCQ MaScQA II 348
Lattice Volume Calculation MCQ Materials Project III 160
Perovskite Stability Prediction MCQ MAST-ML III 480
Diffusion Rate Analysis MCQ Dilute Solute Diffusion Database III 149
L4 Material Safety QA GEN Nature Portfolio III 841
Material toxicity prediction MCQ Toxric III 615
L5 Properties Utilization Analysis GEN Material handbooks I 118
Crystal Structure and Composition Analysis GEN Crystal-LLM III 196
Specified Band Gap Material Generation GEN Material Project III 300
L1 Physics Literature QA MCQ Literature Corpus I 4403
Fundamental Physics Exam MCQ SciQ II 2375
Physics L2 Detailed Understanding MCQ Literature Corpus I 400
Text Summary GEN Literature Corpus I 400
Hypothesis Verification T/F Literature Corpus I 400
Reasoning and Interpretation MCQ Literature Corpus I 400
L3 High School Physics Calculation MCQ tiku.cn II 698
General Physics Calculation MCQ SciEval, SciBench II 800
Physics Formula Derivation MCQ Physics Inference Dataset II 218
L4 Physics Safety QA GEN Nature Portfolio III 342
Laboratory Safety Test MCQ LabExam (ZJU) II 605
L5 Physics Problem Solving GEN Qualifying Exam II 302
@@
Material Project Material-related dataset https://next-gen.materialsproject.org/ OpenSource
Crystal-LLM Crystal-Text dataset https://github.com/facebookresearch/crystal-text-llm OpenSource
MaScQA Material QA dataset https://github.com/M3RG-IITD/MaScQA OpenSource
Nature Portfolio Material literature corpus https://www.nature.com/nature-portfolio CC BY 4.0
MAST-ML Material simulation toolkit https://github.com/uw-cmg/MAST-ML OpenSource
Appendix A5 Prompts for Constructing the Dataset